# Next-to-leading-order QCD corrections to nucleon Dirac form factors

### v2

@@ Abstract
The leading-order perturbative QCD (pQCD) predictions to nucleon electromagnetic form factors were first made in late 70s. Nearly half a century later, we accomplish the calculation of the next-to-leading-order (NLO) QCD corrections to proton and neutron’s Dirac form factors at large momentum transfer, to the leading-twist accuracy in collinear factorization. The NLO hard-scattering kernel is computed in the Krankl and Manashov renormalization scheme. The effect of NLO perturbative corrections turns out to be positive and substantial. Confronting our state-of-the-art pQCD predictions with the available data for nucleon Dirac form factors, in both space-like and time-like regions, imposes a stringent test on the validity of the parameterized forms of nucleon leading-twist light-cone-distribution amplitudes (LCDAs), which have been predicted by a class of QCD sum rules-based models and lattice QCD simulation.
@@
Nucleon electromagnetic form factors (EMFFs), which gauge the distributions of the electric charge and magnetization inside the nucleon, are the fundamental probes to the internal structure of nucleon Gross:2022hyw . The nucleon EMFFs can be directly accessed from the $ep$ elastic scattering experiment, which are encoded in the following nucleon transition matrix element:
$\langle P^{\prime}|j^{\mu}_{\rm em}|P\rangle=\bar{N}\left[F_{1}(Q^{2})\gamma^{\mu}-F_{2}(Q^{2})\frac{\sigma^{\mu\nu}q_{\nu}}{2m_{N}}\right]N,$ (1)
with the electromagnetic current $j^{\mu}_{\rm em}=\sum_{f}e_{f}\bar{\psi}_{f}\gamma^{\mu}\psi_{f}$ ($e_{u}={2\over 3}$, $e_{d}=e_{s}=-{1\over 3})$. $N$, $m_{N}$ denote the Dirac spinor and mass of the nucleon, respectively. $q^{\mu}\equiv P^{\mu}-P^{\prime\mu}$, and $Q^{2}\equiv-q^{2}\geq 0$ characterizes the squared momentum transfer. The scalar functions $F_{1}$, $F_{2}$ are referred to as the (helicity-conserving) Dirac form factor, and the (helicity-flipped) Pauli form factor, respectively.
@@
On the theoretical side, enormous intellectual endeavors have also been invested to unveiling the QCD dynamics beneath the nucleon EMFFs. At small and moderate $Q^{2}$, the measured nucleon EMFFs have been confronted with the theoretical predictions made from various nonperturbative approaches, e.g., vector meson dominance model Iachello:1972nu ; Gari:1984ia ; Lomon:2006xb ; Brodsky:2003gs ; Dubnickova:1992ii ; Chen:2023oqs , constitute quark model Chung:1991st ; Faessler:2005gd ; deMelo:2008rj , cloudy bag model Thomas:1981vc ; Betz:1983dy ; Song:1992ei , Dyson-Schwinger equation Bashir:2012fs ; Cloet:2013gva ; Segovia:2014aza , light-front Hamiltonian Mondal:2019jdg , chiral perturbation theory Schindler:2005ke ; Bauer:2012pv , light-cone sum rule Braun:2001tj , lattice QCD Collins:2011mk ; Alexandrou:2011db ; QCDSF:2017ssq ; Capitani:2015sba ; Alexandrou:2018sjm ; Jang:2019jkn , as well as dispersive analysis Mergell:1995bf ; Belushkin:2006qa .
@@
The lowest-order (LO) hard-scattering kernel for $F_{1}$ was known in the late 70s, shortly after the advent of QCD Lepage:1979za ; Chernyak:1980dj ; Lepage:1980fj ; Chernyak:1977fk . Curiously, notwithstanding nearly half a centaury has elapsed, the next-to-leading order (NLO) perturbative correction to this fundamental observable of nucleon remains unknown until today, obviously due to severe technical obstacles  [^6] Note the status of pQCD predictions for nucleons EMFFs lags far behind that for pion EMFFs, where the one-loop QCD corrections were known in early 80s Field:1981wx ; Dittes:1981aw ; Sarmadi:1982yg ; Braaten:1987yy , and even the two-loop corrections have also become available very recently Chen:2023byr .. The theoretical stagnation for such a basic observable over a very long time is rather unusual, which should be contrasted with the $ep$ deep inelastic scattering, where the three-loop QCD corrections became available about two decades ago Vermaseren:2005qc .
@@
It is the very goal of this work to fill this overdue gap, by accomplishing the calculation of the NLO perturbative correction to the nucleon Dirac form factors, and conducting a comprehensive phenomenological study.
@@
Collinear factorization for nucleon Dirac form factor. At large momentum transfer, the $F_{1}$ can be factorized as the following convolution integrals:
$Q^{4}F_{1}(Q^{2})$ $=$ $\int\!\!\!\!\int[dx][dy]\,\Phi_{3}^{*}(y,\mu_{F})\,\Phi_{3}(x,\mu_{F})$ (2)
×$\times\;T_{H}\left(x,y,Q^{2},\mu_{R}^{2},\mu_{F}^{2}\right)+{\cal O}(1/Q^{2}),$
@@
with the integration measure $[dx]\equiv dx_{1}dx_{2}dx_{3}\delta(1-\sum_{i}x_{i})$. $T_{H}(x,y)$ signifies the hard-scattering kernel, with $\mu_{R}$, $\mu_{F}$ being the renormalization and factorization scales. $\Phi_{3}(x)$ represents the twist-3 nucleon LCDA, $i.e.$, the probability amplitude of finding the three valence quarks inside nucleon carrying the momentum fractions $x_{i}$ ($i=1,2,3$). This factorized structure has been picturised in a cartoon in Fig. 1.
@@
Figure 1: Pictorial illustration of the nucleon EMFF in collinear factorization approach. For concreteness, we show one typical LO and NLO Feynman diagrams for partonic reactions.
@@
For definiteness let us specialize to the proton. The leading-twist-3 LCDA of the proton with helicity $h={1\over 2}$ assumes the following operator definition Chernyak:1983ej ; Braun:2000kw :
[…]
@@
with $P^{2}=m_{N}^{2}$. $i,j,k=1,\cdots,3$ denote the color indices, and $q_{\pm}\equiv{1\over 2}(1\pm\gamma_{5})q$ represents the right and left-handed components of the massless quark field, and $u^{\uparrow}_{p}(P)$ signifies the Dirac spinor of $h={1\over 2}$ proton in relativistic normalization. ${\cal C}\equiv i\gamma^{2}\gamma^{0}$ represents the charge conjugation matrix. $n^{\mu}$ is a light-like vector. For simplicity we have suppressed those light-like gauge links whose role is to ensure gauge invariance. Renormalizing the nonlocal baryonic operator in (3), one recovers the celebrated Efremov-Radyushkin-Brodsky-Lepage (ERBL) evolution equation Lepage:1980fj ; Efremov:1979qk :
$\frac{d\Phi_{3}(x,\mu_{F})}{d\ln\mu^{2}_{F}}=V(x,y)\mathop{\otimes}\limits_{y}\Phi_{3}(y,\mu_{F}),$ (4)
where $\mathop{\otimes}\limits_{y}$ is a shorthand for the convolution integration over $[dy]$.
@@
Perturbative matching in KM scheme. We first sketch how to achieve the desired factorization formula (2). With $Q^{2}\gg\Lambda^{2}_{\rm QCD}$, the transition matrix element in (1) in $d$-dimensional spacetime can be factorized as
@@
$\langle\Pi^{\mu}(Q)\rangle$$\equiv$ $\langle P^{\prime}|j_{\rm em}^{\mu}|P\rangle=\langle P^{\prime}|{\bm{[}}\bar{\cal Q}_{\alpha\beta\gamma}(y){\bm{]}}^{\overline{\rm MS}}|0\rangle$
@@
where $\alpha,\beta,\gamma,\cdots$ denote the spinor indices affiliated with each quark line, ${\cal T}$ signifies the perturbative coefficient function. The three-quark operator ${\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x)$ is defined via the following Fourier transform Anikin:2013aka :
$\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})=\int[dx]e^{-iP_{n}\sum a_{i}x_{i}}{\cal Q}_{\alpha\beta\gamma}(x),$ (6)
where $\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})$ represents the following non-local three-quark operators with light-like separations:
$\widetilde{\cal Q}_{\alpha\beta\gamma}(\{a_{i}\})=\epsilon^{ijk}u^{i}_{\alpha}(a_{1}n)u^{j}_{\beta}(a_{2}n)d^{k}_{\gamma}(a_{3}n),$ (7)
with the light-like gauge links tacitly inserted.
The ${\overline{\rm MS}}$-renormalized three-quark operators in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) is connected with the bare one through
@@
It is advantageous to adopt the renormalization scheme originally proposed by Krankl and Manashov Krankl:2011gch (KM scheme henceforth). The KM scheme has several remarkable virtues. It guarantees the vanishing of evanescent operators Dugan:1990df in $d=4$ dimensions so that one only needs to work on the four-dimensional operator basis. The renormalization procedure in KM scheme preserves Fierz identities between the renormalized operators Anikin:2013aka . The self-consistency of the KM scheme has been checked through the three-loop accuracy Gracey:2012gx . Moreover, this scheme has also been utilized by the lattice QCD practitioners to decipher the baryon LCDA Bali:2015ykx ; RQCD:2019hps ; Bali:2024oxg .
The ${\cal Z}$ factor in (8) can be parameterized in KM scheme as
${\cal Z}_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x,x^{\prime})$ $=$ $\delta_{\alpha\beta\gamma}^{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}\delta(x-x^{\prime})$ (9)
$+\sum_{s=1}{1\over\epsilon^{s}}\,a^{s}_{lmn}\,\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}$
where $\gamma^{(n)}$ is the shorthand for the totally anti-symmetrized product of $n$ $\gamma$-matrices, $\gamma^{(n)}_{\mu_{1}\cdots\mu_{n}}\equiv\gamma^{[\mu_{1}}\gamma^{\mu_{2}}\cdots\gamma^{\mu_{n}]}$. Note that all the Lorentz indices of $\gamma$-matrices are contracted between different quark lines, where there is only one independent way to contract all indices in ${\cal Z}$. For reader’s convenience, the one-loop expression of ${\cal Z}$ is explicitly given in Appendix A.
@@
Note that the hard kernel $T$ is a scalar function, and Lorentz index $\mu$ is affiliated with one of the $\gamma^{(l)}$, $\gamma^{(m)}$ and $\gamma^{(n)}$ matrices.
The central task now is to deduce the coefficient function $T_{lmn}(x,y)$. Since it is insensitive to the long-distance physics, one can replace the incoming and outgoing nucleons by a fictitious nucleon, viz., the free, collinear three-quark states $|q_{1}(u_{1}P)q_{2}(u_{2}P)q_{3}(u_{3}P)\rangle$ and $|q_{1}(v_{1}P)q_{2}(v_{2}P)q_{3}(v_{3}P)\rangle$. As a great simplification of KM scheme, the quark flavor becomes irrelevant in determining ${\cal Z}$ and $T$ Krankl:2011gch .
For a fictitious nucleon, the left-side in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) can be computed in perturbation theory, which can also be expanded in the basis spanned by the totally anti-symmetrized product of $\gamma$-matrices:
$\langle\Pi^{\mu}(Q)\rangle={\cal A}_{lmn}(u,v)\,\bar{u}_{\alpha\beta\gamma}\left[\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}},$
@@
Finally we arrive at the following matching equation to deduce $T$:
${\cal A}_{lmn}(u,v)\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}=\bar{\cal Z}_{\alpha\beta\gamma}^{\alpha_{2}\beta_{2}\gamma_{2}}(y,v)$ (13)
@@
Note that one has to reexpand the products of $\gamma$-marices in the right-hand side, and finally cast them in the basis of the totally anti-symmetrized products of $\gamma$-matrices.
Sketch of the calculation. We start with the partonic reaction $\gamma^{*}+q_{1}(u_{1}P)q_{2}(u_{2}P)q_{3}(u_{3}P)\to q_{1}(v_{1}P^{\prime})q_{2}(v_{2}P^{\prime})q_{3}(v_{3}P^{\prime})$, We use HepLib Feng:2021kha and FeynArts Hahn:2000kx to generate the Feynman diagrams and the corresponding amplitudes. There are about 48 tree-level diagrams, and 2040 one-loop diagrams.
As is evident in Fig. 1, the amplitude for each Feynman diagram entails the following form:
$\bar{u}_{\alpha\beta\gamma}\left[\Gamma^{(1)}_{\alpha\alpha^{\prime}}\otimes\Gamma^{(2)}_{\beta\beta^{\prime}}\otimes\Gamma^{(3)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}},$ (14)
After moving $P\!\!\!\!/$ to the rightmost and $P\!\!\!\!/\,^{\prime}$ to the leftmost repeatedly using the anti-communication relation, and with the aid of Dirac equation, each $\Gamma^{(i)}$ can be expressed in terms of the products of a string of Dirac $\gamma$ matrices, which can be further cast into the linear combination of the totally anti-symmetrized products of $\gamma$ matrices.
The aforementioned operations can be automated by the package HepLib Feng:2021kha . We then use Apart Feng:2012iq and FIRE Smirnov:2019qkx for partial fraction and IBP reduction. We end up with about 100 one-loop two-point and three-point master integrals, which are computed analytically by the differential equation method Kotikov:1990kg ; Remiddi:1997ny . Upon renormalizing the QCD coupling in $\overline{\rm MS}$ scheme, we obtain a UV finite expression of the NLO QCD amplitude ${\cal A}_{lmn}$ in (11), which however still possess uncancelled single infrared pole. From the matching equation (13), we find that, hearteningly, the IR poles in the NLO amplitude are exactly cancelled by the corresponding IR poles in the ${\cal Z}$ factor, so that the short-distance coefficient function $T_{lmn}(x,y)$ is indeed IR finite in a pointwise manner. Therefore it is verified that, the collinear factorization does hold at one-loop order in the baryonic sector.
To recover the form of the standard collinear factorization formula in (2), we substitute the following expression Braun:1999te ; Anikin:2013aka
$\langle 0|{\bm{[}}{\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x){\bm{]}}^{\overline{\rm MS}}|P\rangle=\frac{1}{4}\Big{\{}{\cal V}(P\!\!\!\!/{\cal C})_{\alpha\beta}(\gamma_{5}N)_{\gamma}$ (15)
@@
into (Next-to-leading-order QCD corrections to nucleon Dirac form factors), and use the following relation between ${\cal V}(x),\,{\cal A}(x),\,{\cal T}(x)$ and $\Phi_{3}(x)$:
$\Phi_{3}(x_{1},x_{2},x_{3})={\cal V}(x_{1},x_{2},x_{3})-{\cal A}(x_{1},x_{2},x_{3}),$ (16a)
${\cal T}(x_{1},x_{2},x_{3})=\frac{1}{2}\left[\Phi_{3}(x_{1},x_{3},x_{2})+\Phi_{3}(x_{2},x_{3},x_{1})\right],$
and obtain
$Q^{4}\langle\Pi^{\mu}(Q)\rangle=\Phi_{3}(x)\mathop{\otimes}_{x}T_{H}(x,y)\mathop{\otimes}_{y}\Phi_{3}^{*}(y)\;\bar{N}\gamma^{\mu}N.$ (17)
One then extracts the intended Dirac form factor through NLO accuracy as indicated in (2).
Master formula for nucleon Dirac form factor at NLO. The leading-twist-3 nucleon LCDA is conveniently expanded in the orthogonal polynomials basis Braun:2008ia :
$\Phi_{3}(x,\mu_{F})$ $=$ $f_{\rm N}(\mu_{F})\!\!\sum_{0\leq i\leq j}\!\!a_{ij}(\mu_{F})\phi_{ij}(x)$ (18)
$\phi_{ij}(x)$ $=$ $120x_{1}x_{2}x_{3}\mathcal{P}_{ij}(x),$ (19)
with $f_{\rm N}$ denotes the nucleon decay constant. The first few orthogonal polynomials ${\cal P}_{ij}(x)$ have been tabulated in Braun:2008ia ; Braun:2014wpa (e.g, (12) in Braun:2014wpa ). All the nonperturbative dynamics is encoded in the expansion coefficients $a_{ij}(\mu_{F})$.
For phenomenological usage, it is advantageous to recast (2) from the convolution form into an algebraic one:
@@
with ${\cal H}^{ijkl}$ defined by
${\cal H}^{ijkl}=\phi_{ij}(x)\mathop{\otimes}\limits_{x}T_{H}(x,y,\frac{\mu_{R}^{2}}{Q^{2}},\frac{\mu_{F}^{2}}{Q^{2}})\mathop{\otimes}\limits_{y}\phi_{kl}(y).$ (21)
The short-distance coefficients ${\cal H}^{ijkl}$ can be calculated in perturbation theory. Through NLO accuracy, they can be parameterized as

### v3

@@ Abstract
The leading-order perturbative QCD (pQCD) predictions to nucleon electromagnetic form factors were first made in late 70s. In this Letter for the first time we accomplish the calculation of the next-to-leading-order (NLO) QCD corrections to nucleon’s Dirac form factors at large momentum transfer, to the leading-twist accuracy in collinear factorization approach, specifically within the Krankl and Manashov renormalization scheme. The effect of NLO perturbative corrections turns out to be positive and substantial. Taking the nucleon leading-twist light-cone-distribution amplitudes (LCDAs) determined from the recent lattice simulations as input, we find that the state-of-the-art pQCD predictions significantly underestimate the available nucleon Dirac form factors in both space-like and time-like domains. This nuisance indicates that some additional soft nonfactorizable contribution might be called for to account for the measured nucleon electromagnetic form factor data up to $Q^{2}\approx 30\;{\rm GeV^{2}}$.
@@
Nucleon electromagnetic form factors (EMFFs), which gauge the distributions of the electric charge and magnetization inside the nucleon, are the fundamental probes to the internal structure of nucleon Gross:2022hyw . The nucleon EMFFs can be directly accessed from the $ep$ elastic scattering experiments, which are encoded in the following nucleon transition matrix elements:
$\langle P^{\prime}|j^{\mu}_{\rm em}|P\rangle=\overline{N}\left[F_{1}(Q^{2})\gamma^{\mu}-F_{2}(Q^{2})\frac{i\sigma^{\mu\nu}q_{\nu}}{2m_{N}}\right]N,$ (1)
with the electromagnetic current $j^{\mu}_{\rm em}=\sum_{f}e_{f}\bar{\psi}_{f}\gamma^{\mu}\psi_{f}$. $N$, $m_{N}$ denote the Dirac spinor and mass of the nucleon, respectively. $q^{\mu}\equiv P^{\mu}-P^{\prime\mu}$, and $Q^{2}\equiv-q^{2}\geq 0$ characterizes the squared momentum transfer. The scalar functions $F_{1}$, $F_{2}$ are referred to as the (helicity-conserving) Dirac form factor, and the (helicity-flipped) Pauli form factor, respectively.
@@
On the theoretical side, enormous intellectual endeavors have also been invested to unveiling the QCD dynamics beneath the nucleon EMFFs. At small and moderate $Q^{2}$, the measured nucleon EMFFs have been confronted with the theoretical predictions made from various nonperturbative approaches, e.g., vector meson dominance model Iachello:1972nu ; Gari:1984ia ; Lomon:2006xb ; Brodsky:2003gs ; Dubnickova:1992ii ; Chen:2023oqs , constitute quark model Chung:1991st ; Faessler:2005gd ; deMelo:2008rj , cloudy bag model Thomas:1981vc ; Betz:1983dy ; Song:1992ei , Dyson-Schwinger equation Bashir:2012fs ; Cloet:2013gva ; Segovia:2014aza , light-front Hamiltonian Mondal:2019jdg , chiral perturbation theory Schindler:2005ke ; Bauer:2012pv , light-cone sum rule Passek-Kumericki:2008uqr ; Braun:2001tj , lattice QCD Collins:2011mk ; Alexandrou:2011db ; QCDSF:2017ssq ; Capitani:2015sba ; Alexandrou:2018sjm ; Jang:2019jkn , as well as dispersive analysis Mergell:1995bf ; Belushkin:2006qa .
@@
The lowest-order (LO) hard-scattering kernel for $F_{1}$ was known in the late 70s, shortly after the advent of QCD Lepage:1979za ; Chernyak:1980dj ; Lepage:1980fj ; Chernyak:1977fk . Curiously, notwithstanding nearly half a centaury has elapsed, the next-to-leading order (NLO) perturbative correction to this fundamental observable of nucleon remains elusive until today, obviously due to severe technical obstacles  [^3] Note the status of pQCD predictions for nucleons EMFFs lags far behind that for pion EMFFs, where the one-loop QCD corrections were known in early 80s Field:1981wx ; Dittes:1981aw ; Sarmadi:1982yg ; Braaten:1987yy , and even the two-loop corrections have also become available very recently Chen:2023byr .. The theoretical stagnation for such a basic observable over a very long time is rather unusual, which should be contrasted with the $ep$ deep inelastic scattering, where the three-loop QCD corrections became available about two decades ago Vermaseren:2005qc .
@@
It is the very goal of this Letter to fill this overdue gap, by accomplishing the calculation of the NLO perturbative correction to the nucleon Dirac form factors, and conducting a comprehensive phenomenological investigation.
@@
Collinear factorization for nucleon Dirac form factor. At large momentum transfer, the $F_{1}$ can be factorized as the following convolution integrals  [^4] It is worth mentioning that, this factorization formula is anticipated to be subject to modification starting at the two-loop order, due to emergence of the new leading region contribution Duncan:1979hi ; Kivel:2010ns .
$Q^{4}F_{1}(Q^{2})=\Phi_{3}^{*}(y)\mathop{\otimes}_{x}T_{H}\left(x,y,Q^{2}\right)\mathop{\otimes}_{y}\Phi_{3}(x)+{\cal O}\left(\frac{1}{Q^{2}}\right),$ (2)
@@
where $\mathop{\otimes}\limits_{x}$ is a shorthand for the convolution integration over $dx]\equiv dx_{1}dx_{2}dx_{3}\delta(1-\sum_{i}x_{i})$. $T_{H}(x,y)$ signifies the hard-scattering kernel and $\Phi_{3}(x)$ represents the twist-3 nucleon LCDA. This factorized structure is picturised in a cartoon in Fig. 1.
@@
Figure 1: Pictorial illustration of the nucleon EMFF in collinear factorization. For concreteness, we show one typical LO and NLO Feynman diagrams for partonic reactions.
@@
For definiteness let us specialize to the proton. The leading-twist-3 LCDA of the proton assumes the following operator definition Chernyak:1983ej ; Braun:2000kw :
$\langle 0|\epsilon_{ijk}(u_{+}^{i\,T}(a_{1}n){\cal C}\not\!nu^{j}_{-}(a_{2}n))\not\!nd^{k}_{+}(a_{3}n)|p^{\uparrow}(P)\rangle$
$\quad=-\frac{1}{2}P\cdot n\not\!nu^{\uparrow}_{p}(P)\int[dx]e^{-iP\cdot n\sum x_{i}a_{i}}\Phi_{3}(x),$ (3)
@@
where $i,j,k$ denote the color indices, and $q_{\pm}\equiv{1\over 2}(1\pm\gamma_{5})q$ represents the right and left-handed components of the massless quark field, and $u^{\uparrow}_{p}(P)$ signifies the Dirac spinor of $h={1\over 2}$ proton in relativistic normalization. $\cal C$} represents the charge conjugation matrix and $n^{\mu}$ is a light-like vector. For simplicity we have suppressed those light-like gauge links whose role is to ensure gauge invariance. Renormalizing the nonlocal baryonic operator in (3), one recovers the celebrated Efremov-Radyushkin-Brodsky-Lepage (ERBL) evolution equation Lepage:1980fj ; Efremov:1979qk : ${d\Phi_{3}(x,\mu_{F})}/{d\ln\mu^{2}_{F}}=V(x,y)\mathop{\otimes}\limits_{y}\Phi_{3}(y,\mu_{F}).$
@@
Perturbative matching in KM scheme. We first sketch how to achieve the desired factorization formula (2). With $Q^{2}\gg\Lambda^{2}_{\rm QCD}$, asymptotic freedom indicates that the transition matrix element in (1) in $d$-dimensional spacetime can be factorized as
@@
$\langle\Pi^{\mu}(Q^{2})\rangle$ $\equiv$ $\langle P^{\prime}|j_{\rm em}^{\mu}|P\rangle=\langle P^{\prime}|{\bm{[}}\overline{\cal Q}_{\alpha\beta\gamma}(y){\bm{]}}^{\overline{\rm MS}}|0\rangle$
@@
where $\alpha,\beta,\gamma,\cdots$ denote the spinor indices affiliated with each quark line, $\cal T$} signifies the perturbative coefficient function. The three-quark operator ${\cal Q}_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}(x)$ is defined via the Fourier transform of the non-local three-quark operators with light-like separations $\epsilon^{ijk}u^{i}_{\alpha}(a_{1}n)u^{j}_{\beta}(a_{2}n)d^{k}_{\gamma}(a_{3}n)$, where the light-like gauge links tacitly inserted Anikin:2013aka .
It is advantageous to adopt the renormalization scheme originally proposed by Krankl and Manashov Krankl:2011gch (KM scheme henceforth)  [^5] The KM scheme has several remarkable virtues. It warrants the vanishing of evanescent operators Dugan:1990df in $d=4$ dimensions so that one only needs to work on the four-dimensional operator basis. The renormalization procedure in KM scheme preserves Fierz identities between the renormalized operators Anikin:2013aka . The self-consistency of the KM scheme has been checked through the three-loop accuracy Gracey:2012gx . It is also worth stressing that, the KM scheme has also been routinely utilized by the lattice QCD practitioners to calculate the baryon LCDAs Bali:2015ykx ; RQCD:2019hps ; Bali:2024oxg .. The $\overline{\rm MS}$}-renormalized three-quark operators in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) is connected with the bare one through
@@
For reader’s convenience, the one-loop expression of $\cal Z$} in KM scheme is explicitly given in the supplementary materials Supplemental:Material .
@@
Note that $T_{lmn}$ is a scalar function, and Lorentz index $\mu$ is affiliated with one of the $\gamma^{(l)}$, $\gamma^{(m)}$ and $\gamma^{(n)}$ matrices with $\gamma^{(n)}_{\mu_{1}\cdots\mu_{n}}\equiv\gamma^{[\mu_{1}}\gamma^{\mu_{2}}\cdots\gamma^{\mu_{n}]}$ Krankl:2011gch ; Anikin:2013aka .
The central task now is to deduce the coefficient function $T_{lmn}(x,y)$. Since it is insensitive to the long-distance physics, one can replace the incoming and outgoing nucleons by “fictitious” nucleons, viz., the free, collinear three-quark states $|q_{1}(u_{1}P)q_{2}(u_{2}P)q_{3}(u_{3}P)\rangle$ and $|q_{1}(v_{1}P^{\prime})q_{2}(v_{2}P^{\prime})q_{3}(v_{3}P^{\prime})\rangle$. A virtue of the KM scheme is that the quark flavor structure becomes irrelevant in determining $\cal Z$} and $T$ Krankl:2011gch .
For a fictitious nucleon, the left-side in (Next-to-leading-order QCD corrections to nucleon Dirac form factors) can be computed in perturbation theory:
$\langle\Pi^{\mu}(Q^{2})\rangle={\cal A}_{lmn}(u,v)\,\bar{u}_{\alpha\beta\gamma}\left[\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}},$ (7)
@@
Practically we utilize the following matching equation to deduce $T$ 666 Note that one has to reexpand the products of $\gamma$-marices in the right-hand side of (9), and finally cast them in the basis of the totally anti-symmetrized products of $\gamma$-matrices. :
${\cal A}_{lmn}(u,v)\gamma^{(l)}_{\alpha\alpha^{\prime}}\otimes\gamma^{(m)}_{\beta\beta^{\prime}}\otimes\gamma^{(n)}_{\gamma\gamma^{\prime}}={\overline{\cal Z}}_{\alpha\beta\gamma}^{\alpha_{2}\beta_{2}\gamma_{2}}(y,v)$ (9)
@@
Sketch of the calculation. We start with the partonic reaction $\gamma^{*}+q_{1}q_{2}q_{3}\to q_{1}q_{2}q_{3}$. We use the packages HepLib Feng:2021kha and FeynArts Hahn:2000kx to generate the Feynman diagrams and the corresponding amplitudes. There are about 48 tree-level diagrams, and 2040 one-loop diagrams.
As is evident in Fig. 1, the amplitude in each diagram bears the form $\bar{u}_{\alpha\beta\gamma}\left[\Gamma^{(1)}_{\alpha\alpha^{\prime}}\otimes\Gamma^{(2)}_{\beta\beta^{\prime}}\otimes\Gamma^{(3)}_{\gamma\gamma^{\prime}}\right]u_{\alpha^{\prime}\beta^{\prime}\gamma^{\prime}}$. Moving P/P\!\!\!\!/ to the rightmost and $P\!\!\!\!/\,^{\prime}$ to the leftmost repeatedly using the anti-communication relation  [^7] Note that $\ell\!\!\!/$ also appears ubiquitously in the intermediate stage. By Lorentz covariance, one can trade the loop momentum $\ell$ for the metric tensor or the external momenta $P$ and $P^{\prime}$., and employing Dirac equation, each $\Gamma^{(i)}$ can be expressed as the product of a string of Dirac $\gamma$ matrices, which can be further cast into the form prescribed in (7).
The aforementioned operations can be automated by HepLib Feng:2021kha . We then use Apart Feng:2012iq and FIRE Smirnov:2019qkx for partial fraction and IBP reduction. We end up with about 100 one-loop two-point and three-point master integrals, which are computed using the differential equation method Kotikov:1990kg ; Remiddi:1997ny . Upon renormalizing the QCD coupling in $\overline{\rm MS}$ scheme, we obtain the UV-finite expression of the NLO coefficients ${\cal A}_{lmn}$ in (7), which however still possess uncancelled single infrared pole. Validated by the matching equation (9), we find that, hearteningly, the IR poles in ${\cal A}_{lmn}$ are exactly cancelled by the corresponding poles in $\cal Z$}, so that the short-distance coefficients $T_{lmn}(x,y)$ are indeed IR finite in a pointwise manner. Therefore we verify that the collinear factorization does hold at one-loop order for the nucleon Dirac form factor.
To recover the desired structure in (2), we substitute the following expression Braun:1999te ; Anikin:2013aka
$\hskip-8.5359pt\langle 0|{\bm{[}}{\cal Q}_{\alpha\beta\gamma}(x){\bm{]}}^{\overline{\rm MS}}|P\rangle=\frac{1}{4}\Big{\{}{\cal V}(P\!\!\!\!/{\cal C})_{\alpha\beta}(\gamma_{5}N)_{\gamma}$ (10)
@@
into (Next-to-leading-order QCD corrections to nucleon Dirac form factors). Using the relations among ${\cal V}(x)$, ${\cal A}(x)$, ${\cal T}(x)$ and $\Phi_{3}(x)$ Braun:1999te ; Anikin:2013aka , we obtain $Q^{4}\langle\Pi^{\mu}(Q^{2})\rangle=\Phi_{3}(x)\mathop{\otimes}\limits_{x}T_{H}(x,y)\mathop{\otimes}\limits_{y}\Phi_{3}^{*}(y)\,\bar{N}\gamma^{\mu}N$. The intended Dirac form factor through NLO accuracy can then be predicted in line with (2).
Master formula for nucleon Dirac form factor at NLO. The leading-twist nucleon LCDA is conveniently expanded in the orthogonal polynomials basis Braun:2008ia :
$\Phi_{3}(x,\mu_{F})=f_{\rm N}(\mu_{F})\!\!\sum_{0\leq i\leq j}\!\!a_{ij}(\mu_{F})\phi_{ij}(x),$ (11)
with $\phi_{ij}(x)=120x_{1}x_{2}x_{3}\mathcal{P}_{ij}(x)$ and $f_{\rm N}$ denotes the nucleon decay constant. The first few orthogonal polynomials ${\cal P}_{ij}(x)$ have been tabulated in Braun:2008ia ; Braun:2014wpa (e.g, (12) in Braun:2014wpa ). All the nonperturbative dynamics is encoded in the expansion coefficients $a_{ij}(\mu_{F})$.
For phenomenological usage, it is advantageous to recast (2) from the convolutional form into an algebraic one:
@@
with ${\cal H}^{ijkl}\equiv\phi_{ij}(x)\mathop{\otimes}\limits_{x}T_{H}(x,y)\mathop{\otimes}\limits_{y}\phi_{kl}(y)$.
The short-distance coefficients ${\cal H}^{ijkl}$ can be organized in powers of $\alpha_{s}$. Through NLO accuracy, they can be parameterized as